# Detrimental non-Markovian errors for surface code memory

### v2

@@ I Introduction
The full potential of quantum computers is expected to emerge once fault-tolerant quantum computers have been realized. The main pathway to reach fault tolerance is quantum error correction (QEC) [1, 2, 3], which entails redundantly encoding the quantum information of a few qubits onto the logical subspace of many physical qubits. However, the efficacy of such protocols heavily depends on the underlying noise characteristics affecting the physical system. Crucially, complex, correlated noise can break critical assumptions that quantum error-correcting codes rely on to suppress logical errors [4, 5, 6]. Experimentally, noise affecting QEC protocols shows signs of a rich underlying structure [7, 8, 9], leading to numerous studies focused on mitigating its adverse effects. Beyond spatial inhomogeneity, quantum devices may experience biased noise [10, 11, 12], coherent errors [13, 14], crosstalk [15, 16, 17, 18], noise drift [19, 20], leakage [21, 22, 23, 24], cosmic ray bursts [25, 26, 27, 28], and generally non-Markovian noise [29, 30, 31, 32, 33]. Non-Markovian or temporally correlated noise, in particular, has not been extensively studied in the context of surface codes.
@@
In this paper, we numerically study the performance impact of temporally correlated errors on surface codes—a prominent family of quantum error-correcting codes that are key to many strategies for achieving practical, fault-tolerant quantum computing [34, 35, 36]. Key to our analysis is comparing surface code performance under correlated noise versus independent noise while keeping marginal error rates of each qubit constant, thereby isolating the impact of the structure of noise rather than the overall error rate. Our models explore a range of temporally correlated error patterns, from pairwise to multi-time correlations, and polynomially decaying to exponentially decaying correlations. We find that certain temporally correlated noise structures are catastrophic to logical error rate scaling in surface code performance. Specifically, multi-time streaks of correlated errors affecting syndrome qubits, or two-qubit entangling gates, can result in strongly sub-exponential logical error rate suppression in quantum memory experiments. These findings highlight the need for deeper understanding and mitigation of such correlated errors in the pursuit to build a fault-tolerant quantum computer.
@@ II.1 Surface codes and circuit-level noise models
$\mathcal{E}^{1}_{X}(p)=(1-p)\rho+pX\rho X$ (1a)
$\mathcal{E}^{1}_{D}(p)=(1-p)\rho+\frac{p}{3}(X\rho X+Y\rho Y+Z\rho Z)$ (1b)
$\mathcal{E}^{2}_{D}(p)=(1-p)\rho+\frac{p}{15}\sum_{P\in\{I,X,Y,Z\}^{\otimes 2}\setminus\{II\}}P\rho P.$ (1c)
In standard circuit-level noise, $p$, termed the characteristic physical error rate, is identical for each channel. Note that in some variations of standard circuit-level noise, Class 0 errors are concatenated to a single identity gate followed by a depolarizing error channel at the beginning of the round, which will be the convention adopted in this paper.
@@ II.2 Detector error models
where $\oplus$ represents binary addition. Detector $D_{s,t}$ is effectively a parity operation on $M_{s,t}$ at times $t-1$ and $t$. Note that $M_{s,t}$ is only explicitly measured at times $t\in\{1,2,...,n\}$. $M_{s,0}$ is inferred directly from the data qubit state preparation step, and $M_{s,n+1}$ is inferred from the final data qubit measurement outcomes, defined as $M_{s,n+1}=\bigoplus_{u\in N(s)}M_{u,n+1}$, where $M_{u,n+1}$ is the final measurement outcome of a neighboring data qubit at spatial coordinates $u\in N(s)$. Furthermore, for a logical $Z$ basis memory experiment, detection events are not defined at $X$ syndromes at $t=1$ and $t=n+1$ (and vice versa for $X$ basis memory experiments), resulting in $\frac{m}{2}+(n-1)m+\frac{m}{2}=n\times m$ total number of detectors.
@@ III.1 Noise models
Figure 3: Logical error rate per round versus surface code distance under temporally correlated circuit-level noise models. Error bars correspond to 95% confidence intervals (up to marker size). Memory experiments are $2d$ rounds long and use minimum weight perfect matching to decode logical errors. Under the pairwise correlated noise model (top), error pairs occur with probability $Ap/\Delta t^{2}$, where $\Delta t=|t_{1}-t_{2}|$ is the time separation in rounds. Under the streaky correlated noise model (bottom), error streaks occur with probability $Ap/t^{2}$, where $t$ is the streak length in rounds. In both models, $A_{\text{Class 0}}=A_{\text{Class 1}}=2A_{\text{Class 2}}$. In each plot, we compare performance under correlated noise model with performance under the marginalized independent model, where errors are temporally uncorrelated, but have the same marginal error rates as the correlated model.
Figure 4: Threshold plots under temporally correlated circuit-level noise (pairwise and streaky) and marginalized independent noise. The point at which trendlines intersect, known as the accuracy threshold, represents the physical error rate below which increasing the code distance reduces the logical error rate. For both the pairwise correlated and marginalized independent model, we observe similar thresholds of approximately 0.6%. Additionally, both models exhibit comparable scaling behavior between logical and physical error rates, as previously illustrated in Fig. 3. In contrast, for the streaky correlated and marginalized independent model, with respective thresholds of 0.3% and 0.4%, we observe sub-exponential suppression of logical error rates in the streaky correlated model.
@@
We consider two types of temporally correlated error structures. The first structure involves pairwise correlations, which introduce two-time correlated errors on the same qubit (or same pair of qubits in the case of Class 2 errors) at two different rounds. The two-time error channel depends on the class; for correlated Class 0 errors, a qubit experiences a two-time depolarizing channel at times $t,t^{\prime}$ with errors denoted by $I_{t}X_{t^{\prime}},I_{t}Y_{t^{\prime}},I_{t}Z_{t^{\prime}},...,Z_{t}Z_{t^{\prime}}$ each occurring with equal $1/15$ probability (this is the two-time equivalent to the two-qubit depolarizing channel shown in Eq. 1c). This model can be viewed as the temporal analog of the spatially correlated nonlocal two-qubit error model in Ref. [4]. For correlated Class 1 errors, a qubit experiences a two-time $X_{t}X_{t^{\prime}}$ error, and for correlated Class 2 errors, a pair of qubits experiences a two-time, two-qubit depolarizing error after a CNOT operation. Every round pair experiences a probability of a correlated pairwise error, equating to $\binom{n}{2}=n(n-1)/2$ probability rolls per qubit (since each qubit independently experiences temporally correlated errors), where $n$ is the number of rounds. The probability of any pair of errors decays polynomially or exponentially with round separation. In the polynomial decay model, a pair of errors occur with probability $Ap/\Delta t^{n}$, where $\Delta t=|t_{1}-t_{2}|$ is the distance in rounds, $p$ is the characteristic error rate and $A,n$ are model parameters. Similarly, in the exponential decay model, a pair of errors occur with probability $Ap/n^{\Delta t}$.
@@ III.2 Numerical simulations
Elements of $\bm{M}$ are populated according to a predefined temporally correlated noise model. For example, in the Class 1 temporally correlated pairwise and streaky model, we first construct a probability matrix $\bm{R}=R_{ik}$, where each element $R_{ik}$ gives the probability of a correlated error affecting qubit $i$ at times $t_{1}$ and $t_{2}$, denoted by $k$. This probability is calculated according to the correlation decay model, either polynomial ($Ap/\Delta t^{n}$) or exponential ($Ap/n^{\Delta t}$), as defined in Section III.1. Sampling $\bm{R}$ yields a binary outcome matrix $\bm{O}=O_{ik}$, indicating when and where correlated error events have occurred. These events in $\bm{O}$ are then mapped to individual errors in $\bm{M}$ via a transformation matrix $\bm{T}=T_{kj}$, which yields error mask elements $M_{ij}$ as
@@
$M_{ij}=\bigvee_{k}O_{ik}T_{kj}.$ (4)
Full details are contained in the source code for this project [66].
To decode errors, we employ minimum weight perfect matching via PyMatching [67], inputting the same marginalized independent detector error model for both sets of syndrome data. Each memory experiment is $2d$ rounds long (or two $d\times d\times d$ blocks). We implement polynomially decaying correlations for both the pairwise and streaky correlated model, parameterized by $A_{\text{Class 0}}=A_{\text{Class 1}}=2A_{\text{Class 2}}=1$, and $n=2$. We set the characteristic physical error rate $p=1\times 10^{-3}$, or approximately a tenth of the threshold under standard circuit-level noise.
@@ III.3 Results
In sharp contrast, we observe sub-exponential suppression of logical error rate with increasing code distance, for the temporally correlated streak model from distance 3 to distance 15. We find no realistic distance projection to reach teraquop logical error rates. It is important to reiterate that the marginal per qubit error rates (with characteristic physical error rate $p=10^{-3}$) are well under threshold, as demonstrated by surface code performance under the marginalized independent model (Fig. 3). Our results show that sufficiently strong multi-time correlations in circuit-level noise can be detrimental to achieving fault tolerance with the surface code. Specifically, streaky correlations decaying quadratically or slower will lead to sub-exponential logical error rate suppression. At distance 15, the logical per round error rate under the correlated model is measured to be $2.059\cdot 10^{-5}$, approximately 58 times the error rate under the marginalized independent model of $3.566\cdot 10^{-7}$. Under the marginalized independent model, we fit a trend of $p_{L}=2.51\cdot 10^{-3}e^{-5.95\cdot 10^{-1}d}$ and estimate a teraquop distance projection of $37$.
Figure 4 presents threshold plots for temporally correlated noise models compared to independent noise models with the same marginal error rates. Both the streaky correlated model and its corresponding independent model show similar thresholds in the 0.3%–0.4% range. However, while the independent model exhibits exponential suppression of the logical error rate with increasing code distance below the accuracy threshold, the correlated model follows a power-law suppression. This difference can result in impractically high qubit overheads when aiming for logical error rates in the teraquop regime. In summary, these results so far indicate extra strategies may be required to address circuit-level noise exhibiting multi-time correlations in surface codes. In the next section, we further breakdown the cause of sub-optimal surface code performance under temporally correlated noise by isolating time correlations to single classes of errors.
@@ IV.1 Surface code distance projections by correlated error class
Figure 5: Surface code memory logical error rate versus code distance for various noise models where correlated errors are restricted to a single error class. Error bars represent 95% confidence intervals. Memory experiments are $2d$ rounds long, and errors are decoded using minimum weight perfect matching. All models follow the polynomial decay model $Ap/\Delta t^{n}$ with parameters $A=1,n=2,p=2\times 10^{-3}$ (i.e. quadratically decaying correlations) except for the correlated Class 2 models where $A=0.5$. Models experience independent noise beyond their correlated error class, e.g., the correlated Class 0 models additionally experience uncorrelated Class 1 and Class 2 errors. Correlated results are plotted against independent results, where the marginal error rates are fixed at each qubit at each round.
In our first analysis, we compare the performance of the surface code under various temporally correlated noise models where correlations are restricted to a single error class. We run 10 million trials of stabilizer simulations of surface code memory experiments with $2d$ rounds, decoded using minimum weight perfect matching, inputting the same detector error model for both the correlated models and the completely independent model. All models experience independent circuit-level noise beyond their specific correlated error class. For example, the model with temporally correlated Class 0 errors will additionally experience uncorrelated Class 1 and Class 2 (but not Class 0) errors described by the characteristic physical error rate p. Specifically, this involves an $X$ error occurring with probability $p$ before measurement and after reset, and a $p/15$ probability each of an $IX,IY,IZ,...,ZZ$ error occurring after each two-qubit gate. Both pairwise and streaky correlated noise models follow the polynomial decay model $Ap/t^{n}$ with parameters $A=1,n=2,p=2\times 10^{-3}$, except for the correlated Class 2 models where $A=0.5$. We compare correlated models with independent error models, fixing the marginal error rates for each qubit at each timestep.
@@
The temporally correlated Class 2 noise models exhibit similar behavior. The correlated pairwise model and its corresponding independent model display similar trends of $p_{L}=9.71\cdot 10^{-3}e^{-6.49\cdot 10^{-1}d}$ and $p_{L}=6.66\cdot 10^{-3}e^{-6.14\cdot 10^{-1}d}$, with distance projections $36$ and $37$, respectively. For the correlated Class 2 noise model, we find sub-exponential suppression of logical error rate with increasing distance, with no reasonable teraquop distance projection. The surface code thus does not favorably scale with less than quadratically decaying temporally correlated Class 2 streak errors. For the marginalized independent model, we measure a trend of $p_{L}=3.58\cdot 10^{-3}e^{-3.75\cdot 10^{-1}d}$, and a fault-tolerant distance projection of $59$. The relatively high distance projections in Class 2 models are attributed to the surface code’s increased sensitivity to Class 2 error rates.

### v3

@@ I Introduction
The full potential of quantum computers is expected to emerge once fault-tolerant quantum computers have been realized. The main pathway to reach fault tolerance is quantum error correction (QEC) [1, 2, 3], which entails redundantly encoding the quantum information of a few qubits onto the logical subspace of many physical qubits. However, the efficacy of such protocols heavily depends on the underlying noise characteristics affecting the physical system. Crucially, complex, correlated noise can break critical assumptions that quantum error-correcting codes rely on to suppress logical errors [4, 5, 6, 7]. Experimentally, noise affecting QEC protocols shows signs of a rich underlying structure [8, 9, 10], leading to numerous studies focused on mitigating its adverse effects. Beyond spatial inhomogeneity, quantum devices may experience biased noise [11, 12, 13], coherent errors [14, 15], crosstalk [16, 17, 18, 19], noise drift [20, 21], leakage [22, 23, 24, 25], cosmic ray bursts [26, 27, 28, 29], and generally non-Markovian noise [30, 31, 32, 33, 34]. Non-Markovian or temporally correlated noise, in particular, has not been extensively studied in the context of surface codes.
@@
In this paper, we numerically study the performance impact of temporally correlated errors on surface codes—a prominent family of quantum error-correcting codes that are key to many strategies for achieving practical, fault-tolerant quantum computing [35, 36, 37, 38]. Key to our analysis is comparing surface code performance under correlated noise versus independent noise while keeping marginal error rates of each qubit constant, thereby isolating the impact of the structure of noise rather than the overall error rate. Our models explore a range of temporally correlated error patterns, from pairwise to multi-time correlations, and polynomially decaying to exponentially decaying correlations. We find that certain temporally correlated noise structures are catastrophic to logical error rate scaling in surface code performance. Specifically, multi-time streaks of correlated errors affecting syndrome qubits, or two-qubit entangling gates, can result in strongly slower-than-exponential logical error rate suppression in quantum memory experiments. These findings highlight the need for deeper understanding and mitigation of such correlated errors in the pursuit to build a fault-tolerant quantum computer.
@@ II.1 Surface codes and circuit-level noise models
$\mathcal{E}^{1}_{X}(p{,\rho})=(1-p)\rho+pX\rho X$ (1a)
$\mathcal{E}^{1}_{D}(p{,\rho})=(1-p)\rho+\frac{p}{3}(X\rho X+Y\rho Y+Z\rho Z)$ (1b)
$\mathcal{E}^{2}_{D}(p{,\rho})=(1-p)\rho+\frac{p}{15}\sum_{P\in\{I,X,Y,Z\}^{\otimes 2}\setminus\{II\}}P\rho P.$ (1c)
In standard circuit-level noise $p$ is identical for each channel. Note that in some variations of standard circuit-level noise, Class 0 errors are concatenated to a single identity gate followed by a depolarizing error channel at the beginning of the round, which will be the convention adopted in this paper.
@@ II.2 Detector error models
where $\oplus$ represents modulo 2 addition. Detector $D_{s,t}$ is effectively a parity operation on $M_{s,t}$ at times $t-1$ and $t$. Note that $M_{s,t}$ is only explicitly measured at times $t\in\{1,2,...,N\}$. $M_{s,0}$ is inferred directly from the data qubit state preparation step, and $M_{s,N+1}$ is inferred from the final data qubit measurement outcomes, defined as $M_{s,N+1}=\bigoplus_{u\in G(s)}M_{u,N+1}$, where $M_{u,N+1}$ is the final measurement outcome of a neighboring data qubit at spatial coordinates $u\in G(s)$. Furthermore, for a logical $Z$ basis memory experiment, detection events are not defined at $X$ syndromes at $t=1$ and $t=n+1$ (and vice versa for $X$ basis memory experiments), resulting in $\frac{m}{2}+(N-1)m+\frac{m}{2}=N\times m$ total number of detectors.
@@ III.1 Noise models
Figure 3: Logical error rate per round versus surface code distance under temporally correlated circuit-level noise models. Error bars correspond to 95% confidence intervals (up to marker size). Memory experiments are $2d$ rounds long and use minimum weight perfect matching to decode logical errors. Under the pairwise correlated noise model (top), error pairs occur with probability $Aq/\Delta t^{2}$, where $\Delta t=|t_{1}-t_{2}|$ is the time separation in rounds. Under the streaky correlated noise model (bottom), error streaks occur with probability $Aq/t^{2}$, where $t$ is the streak length in rounds. In both models, $A_{\text{Class 0}}=A_{\text{Class 1}}=2A_{\text{Class 2}}$. In each plot, we compare performance under correlated noise model with performance under the marginalized independent model, where errors are temporally uncorrelated, but have the same marginal error rates as the correlated model.
Figure 4: Threshold plots under temporally correlated circuit-level noise (pairwise and streaky) and marginalized independent noise. The point at which trendlines intersect, known as the accuracy threshold, represents the physical error rate below which increasing the code distance reduces the logical error rate. For both the pairwise correlated and marginalized independent model, we observe similar thresholds of approximately 0.6%. Additionally, both models exhibit comparable scaling behavior between logical and physical error rates, as previously illustrated in Fig. 3. In contrast, for the streaky correlated and marginalized independent model, with respective apparent thresholds of 0.3% and 0.4%, we observe slower-than-exponential suppression of logical error rates in the streaky correlated model up to distance 13.
@@
We consider two types of temporally correlated error structures. The first structure involves pairwise correlations, which introduce two-time correlated errors on the same qubit (or same pair of qubits in the case of Class 2 errors) at two different rounds. The two-time error channel depends on the class; for correlated Class 0 errors, a qubit experiences a two-time depolarizing channel at times $t,t^{\prime}$ with errors denoted by $I_{t}X_{t^{\prime}},I_{t}Y_{t^{\prime}},I_{t}Z_{t^{\prime}},...,Z_{t}Z_{t^{\prime}}$ each occurring with equal $1/15$ probability (this is the two-time equivalent to the two-qubit depolarizing channel shown in Eq. 1c). This model can be viewed as the temporal analog of the spatially correlated nonlocal two-qubit error model in Ref. [5]. For correlated Class 1 errors, a qubit experiences a two-time $X_{t}X_{t^{\prime}}$ error, and for correlated Class 2 errors, a pair of qubits experiences a two-time, two-qubit depolarizing error after a CNOT operation. Every round pair experiences a probability of a correlated pairwise error, equating to $\binom{N}{2}=N(N-1)/2$ probability rolls per qubit (since each qubit independently experiences temporally correlated errors), where $N$ is the number of rounds. The probability of any pair of errors decays polynomially or exponentially with round separation. In the polynomial decay model, a pair of errors occur with probability $Aq/\Delta t^{n}$, where $\Delta t=|t_{1}-t_{2}|$ is the distance in rounds, $q$ is the characteristic correlated error parameter and $A,n$ are model parameters. Similarly, in the exponential decay model, a pair of errors occur with probability $Aq/n^{\Delta t}$.
@@ III.2 Numerical simulations
Elements of $\bm{M}$ are populated according to a predefined temporally correlated noise model. For example, in the Class 1 temporally correlated pairwise and streaky model, we first construct a probability matrix $\bm{R}=R_{ik}$, where each element $R_{ik}$ gives the probability of a correlated error affecting qubit $i$ at times $t_{1}$ and $t_{2}$, denoted by $k$. This probability is calculated according to the correlation decay model, either polynomial ($Aq/\Delta t^{n}$) or exponential ($Aq/n^{\Delta t}$), as defined in Section III.1. Sampling $\bm{R}$ yields a binary outcome matrix $\bm{O}=O_{ik}$, indicating when and where correlated error events have occurred. These events in $\bm{O}$ are then mapped to individual errors in $\bm{M}$ via a transformation matrix $\bm{T}=T_{kj}$, which yields error mask elements $M_{ij}$ as
@@
$M_{ij}=\bigvee_{k}O_{ik}T_{kj},$ (4)
where $\bigvee$ is the logical OR operation. Full details are contained in the source code for this project [67].
To decode errors, we employ minimum weight perfect matching via PyMatching [68], inputting the same marginalized independent detector error model for both sets of syndrome data. Each memory experiment is $2d$ rounds long (or two $d\times d\times d$ blocks). This choice is primarily practical, balancing computational constraints with the need to capture temporal correlations in sufficiently long memory experiments. In realistic fault-tolerant quantum computations, logical qubits often idle for many $d$ rounds, and these correlations are expected to persist or may become even more pronounced in longer experiments. We implement polynomially decaying correlations for both the pairwise and streaky correlated model, parameterized by $A_{\text{Class 0}}=A_{\text{Class 1}}=2A_{\text{Class 2}}=1$, and $n=2$. We set the characteristic correlated error parameter $q=1\times 10^{-3}$.
@@ III.3 Results
In sharp contrast, the temporally correlated streak model exhibits slower-than-exponential suppression of the logical error rate with increasing code distance (from distance 3 to 15). Even when fitting a power-law decay as a best-case approximation and extrapolating, we find no realistic distance projection to achieve teraquop logical error rates. Critically, this projection assumes continued monotonic suppression of logical errors, disregarding alternative asymptotic behaviors such as saturation or a non-monotonic increase in the logical error rate. We have not definitively ruled out these more detrimental scenarios. It is important to reiterate that the marginal per qubit error rates are well under threshold, as demonstrated by surface code performance under the marginalized independent model (Fig. 3). Our results show that sufficiently strong multi-time correlations in circuit-level noise can be detrimental to achieving fault tolerance with the surface code. Specifically, streaky correlations decaying quadratically or slower will, at best, result in slower-than-exponential logical error rate suppression. At distance 15, the logical per round error rate under the correlated model is measured to be $2.059\cdot 10^{-5}$, approximately 58 times the error rate under the marginalized independent model of $3.566\cdot 10^{-7}$. Under the marginalized independent model, we fit a trend of $p_{L}=2.51\cdot 10^{-3}e^{-5.95\cdot 10^{-1}d}$ and estimate a teraquop distance projection of $37$.
Figure 4 presents threshold plots for temporally correlated noise models compared to independent noise models with the same marginal error rates. Both the streaky correlated model and its corresponding independent model show similar apparent thresholds in the 0.3%–0.4% range. The term apparent threshold refers to the initial intersection point of logical error rate trends, but it may not necessarily correspond to a true threshold as conventionally defined (exponential suppression as a function of code distance for surface codes). In fact, for the streaky correlated model, there are strong indications that a true threshold does not exist under a standard decoder. As we demonstrate in Appendix B, further simulations reveal that for sufficiently high physical error rates, the logical error rate is a non-monotonic function of the code distance, eventually increasing after an initial decrease. While the independent model achieves exponential suppression of the logical error rate with increasing code distance below the accuracy threshold, the correlated model exhibits significantly weaker suppression. As a result, reaching teraquop-level logical error rates may require impractically large qubit overheads—or may be impossible if the logical error rate suppression follows a more pathological trend. In summary, these results so far indicate extra strategies may be required to address circuit-level noise exhibiting multi-time correlations in surface codes. In the next section, we further breakdown the cause of suboptimal surface code performance under temporally correlated noise by isolating time correlations to single classes of errors.
@@ IV.1 Surface code distance projections by correlated error class
Figure 5: Surface code memory logical error rate versus code distance for various noise models where correlated errors are restricted to a single error class. Error bars represent 95% confidence intervals. Memory experiments are $2d$ rounds long, and errors are decoded using minimum weight perfect matching. All models follow the polynomial decay model $Aq/\Delta t^{n}$ with parameters $A=1,n=2,q=2\times 10^{-3}$ (i.e. quadratically decaying correlations) except for the correlated Class 2 models where $A=0.5$. Models experience independent noise beyond their correlated error class, e.g., the correlated Class 0 models additionally experience uncorrelated Class 1 and Class 2 errors. Correlated results are plotted against independent results, where the marginal error rates are fixed at each qubit at each round.
In our first analysis, we compare the performance of the surface code under various temporally correlated noise models where correlations are restricted to a single error class. We run 10 million trials of stabilizer simulations of surface code memory experiments with $2d$ rounds, decoded using minimum weight perfect matching, inputting the same detector error model for both the correlated models and the completely independent model. All models experience independent circuit-level noise beyond their specific correlated error class. For example, the model with temporally correlated Class 0 errors will additionally experience uncorrelated Class 1 and Class 2 (but not Class 0) errors occur with probability $p=q$, the characteristic correlated error parameter. Specifically, this involves an $X$ error occurring with probability $p$ before measurement and after reset, and a $p/15$ probability each of an $IX,IY,IZ,...,ZZ$ error occurring after each two-qubit gate. Both pairwise and streaky correlated noise models follow the polynomial decay model $Aq/t^{n}$ with parameters $A=1,n=2,q=2\times 10^{-3}$, except for the correlated Class 2 models where $A=0.5$. We compare correlated models with independent error models, fixing the marginal error rates for each qubit at each timestep.
@@
The temporally correlated Class 2 noise models exhibit similar behavior. The correlated pairwise model and its corresponding independent model display similar trends of $p_{L}=9.71\cdot 10^{-3}e^{-6.49\cdot 10^{-1}d}$ and $p_{L}=6.66\cdot 10^{-3}e^{-6.14\cdot 10^{-1}d}$, with distance projections $36$ and $37$, respectively. For the correlated Class 2 noise model, we find slower-than-exponential suppression of logical error rate with increasing distance, with no reasonable teraquop distance projection. The surface code thus does not favorably scale with less than quadratically decaying temporally correlated Class 2 streak errors. For the marginalized independent model, we measure a trend of $p_{L}=3.58\cdot 10^{-3}e^{-3.75\cdot 10^{-1}d}$, and a fault-tolerant distance projection of $59$. The relatively high distance projections in Class 2 models are attributed to the surface code’s increased sensitivity to Class 2 error rates.